9.) A 50 ml solution of unknown base contains 0.0012 mols of base (2.94 g/L) in solution.
Calculate the molar mass of the solution, and use that to calculate the concentration and
pH of the solution.
10.) Which of the 4 bases was the base from problem 9?
a) KOH
b) NaNH2
c) NH3
d) NaClO4
1
Expert's answer
2020-05-10T14:58:11-0400

9) Find the mass of the unknown base:


m(base) = V*ρi/1000 = 50*2.95/1000 = 0.147 g


The molar mass of the solute = m(base)/n(base) = 0.147/0.0012 = 122.5 g/mol


Molar concentration of solution = 0.0012/0.050 = 0.024 mol/L


10) Calculate molar mass for following compounds:


a) M(KOH) = 56 g/mol

b) M(NaNH2) = 39 g/mol

c) M(NH3) = 17 g/mol

d) M(NaClO4) = 122.5 g/mol


As we can see, the unknown base is (d) NaClO4. NaClO4 solution has a neutral pH, the salt is formed by a strong base and a strong acid.


Answer: molar mass of the solute = 122.5 g/mol and molar concentration of solution = 0.024 mol/L. Unknown base is (d) NaClO4 and solution has a neutral pH (pH = 7)